OpenStudy (anonymous):

A salesperson is paid a flat rate plus a fixed percentage of her sales. Last month, she sold $16,000 worth of goods and was paid $1,600. Two months ago, she had $12,000 in sales and was paid $1,400. This month she sold $11,000 worth of goods. How much will she be paid?

undeadknight26 (undeadknight26):

f + 16000p = 1600 f + 12000p = 1400

undeadknight26 (undeadknight26):

f = 800 p = 0.05 (or 5%)

undeadknight26 (undeadknight26):

800 + 0.05 * 11,000 = $1,350

undeadknight26 (undeadknight26):

Hope i helped?

OpenStudy (anonymous):

thank you